# Pre-trained Universal Medical Image Transformer

### 12 Dec 2023 v1

@@ Introduction
We specifically employ sum pooling for weight reduction to maintain the mean and variance of the output, similar to the weight inflation technique used to adapt pre-trained weights from 2D models to 3D models (Zhang et al. 2022).
Universal Visual Tokenizer
A convolution-based universal visual tokenizer for medical images can be constructed with the proposed SAC module. The encoder of the tokenizer resembles a typical convolutional backbone, but with each $3^{3}$ convolution and downsampling convolution replaced by the corresponding variant of the SAC module. Given an input image with voxel spacing information, it outputs a feature map with $2^{-4}$ of the original resolution on the in-plane dimensions and $2^{-\max\{4-\mathrm{DA},0\}}$ of the original resolution on the depth dimension. The decoder is composed of a stack of the $3^{3}$ convolution and the upsampling variants of the SAC module. It gradually upsamples the quantized feature map output by the encoder, and finally outputs a reconstruction of the original image.
Universal ViT
The patch embedding layer of a ViT divides the input image into a sequence of non-overlapping patches, and maps each patch to an embedding vector through a linear projection. This process is equivalent to processing the input image with a downsampling convolution that has identical kernel size and stride, and flattening the spatial dimensions of the convolution output. Therefore, the downsampling convolution can be replaced by the downsampling variant of the SAC module to accommodate spatially heterogeneous input images. Furthermore, the transformer blocks following the patch embedding layer in ViT are inherently able to process sequences with variable lengths. As a result, a ViT enhanced with the proposed SAC module in the patch embedding layer becomes a universal backbone capable of handling a wide range of medical images with various imaging modalities and spatial properties.
Figure 1: Illustration of our proposed method for pre-training a universal visual tokenizer and a universal ViT for medical images. Our model can process a wide range of images with a unified model structure based on the proposed SAC module.
Visual Tokenizer Details
Existing pre-trained visual tokenizers can suffer from issues of severe codebook collapse, low codebook utilization, and perturbed reconstruction process introduced by stochastic quantization (Zhang et al. 2023). We propose several enhancements to make the visual tokenizer produce more robust tokenization output.
@@
Extended Prior Distribution Regularization
During experiments, we find that using the soft token representation alone still results in low codebook utilization. The prior distribution regularization (Zhang et al. 2023) provides a possible direction of solution. It calculates a distribution $p_{\mathrm{post}}$ by averaging the one-hot distributions over the codebook of all sampled discrete tokens, and minimizes $D_{\mathrm{KL}}\left(p_{\mathrm{post}}\parallel p_{\mathrm{prior}}\right)$, where $p_{\mathrm{prior}}$ is a uniform distribution. This technique can be effortlessly generalized to the soft token representation by replacing the one-hot distributions with the general categorical distributions of tokens.
In fact, there exists a constructive interpretation when using this technique with the soft token representation. Suppose the encoder outputs a $D\times H\times W$ grid of token distributions, let $G_{D,H,W}$ denote the set of discrete coordinates of all cells within the grid. For each position $s\in G_{D,H,W}$ on the grid, we define a random variable $I_{s}$ denoting the index of the token in the codebook at $s$ and its distribution to be $q^{(s)}$. We define another random variable $S$ representing a random position on the grid that is uniformly distributed over $G_{D,H,W}$. It can be shown that the averaged distribution of all token distributions on the grid is exactly the distribution of $I_{S}$:
$=\frac{1}{DHW}\sum_{s\in G_{D,H,W}}\mathrm{P}(I_{s}=i).$ (3)
Intuitively, when the distribution of $I_{S}$ is close to the uniform distribution, individual tokens occur at a randomly sampled position with close probabilities. This benefits the diversity of the learned distributions.
On the other hand, to prevent all distributions collapsing to the uniform distribution, we introduce another objective maximizing $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ to increase the sharpness of the learned distributions. Since $p_{\mathrm{prior}}$ is a uniform distribution, it can be shown that for any distribution $p$, $D_{\mathrm{KL}}\left(p\parallel p_{\mathrm{prior}}\right)=-\mathrm{H}\left(p\right)+\ln\left\lvert\mathcal{V}\right\rvert$, where $\mathrm{H}\left(p\right)=-\sum_{i=1}^{|\mathcal{V}|}p_{i}\ln p_{i}$. Therefore, the extended prior distribution regularization can be implemented as minimizing the following objective:
@@
where $\lambda_{1},\lambda_{2}$ are positive hyper-parameters.
Optimization Objective
Besides the basic L1 reconstruction loss $L_{\mathrm{rec}}$, the training objective of our visual tokenizer includes a perceptual loss $L_{\mathrm{percep}}$ (Zhang et al. 2018) and a discriminative loss $L_{\mathrm{disc}}$ by adversarily training a patch-based distriminator $P$ following VQGAN, and the aforementioned extended prior distribution regularization $L_{\mathrm{reg}}$. During our experiments, we find that using the gradient-based adaptive weight for $L_{\mathrm{disc}}$ proposed in VQGAN results in failure of convergence. Instead, we set the weight for $L_{\mathrm{disc}}$ as a fixed hyper-parameter $\lambda$, and include $L_{\mathrm{disc}}$ in the total objective only when the loss function value of the discriminator $L_{P}$ is smaller than a given threshold $t_{\mathrm{disc}}$. In conclusion, the final optimization objective is minimizing:
$L_{\mathrm{VT}}=L_{\mathrm{rec}}+L_{\mathrm{percep}}+[L_{P}<t_{\mathrm{disc}}]L_{\mathrm{disc}}+L_{\mathrm{reg}},$ (5)
where the []delimited-[][$\cdot$][ ⋅ ] operator is Iverson bracket. The overall structure of our universal visual tokenizer is illustrated in the upper part of Figure 1.
Transformer Architecture
Our ViT model employs a plain ViT macro architecture that is composed of a patch embedding layer and a stack of transformer blocks (Dosovitskiy et al. 2021). The micro architecture of the transformer block follows EVA-02 (Fang et al. 2023a), which incorporates several advancements in language models for visual representation learning, including the 2D rotary position embedding (RoPE) (Su et al. 2021). In short, for each $d$-dimensional query or key vector $\bm{a}=(a_{0},a_{1},\dots,a_{d-1})$ in an attention block, 1D RoPE rotates the pair $(a_{2i},a_{2i+1})$ on the 2D plane by a radian of $\omega_{i}t$, where $t$ is the position index, and it is commonly chosen that $\omega_{i}=b^{-\frac{2i}{d}}$ with a base $b$. For a given position on a 2D image coordinated with $(t_{x},t_{y})$, 2D RoPE extends 1D RoPE by rotating the first half of the vector according to $t_{x}$ and the second half according to $t_{y}$.
We further extend 2D RoPE to work with 3D medical images for any 3D coordinate $(t_{x},t_{y},t_{z})$ by adding an additional rotation radian. Specifically, let $\theta_{i}$ be the rotation radian given by the 2D RoPE at $(t_{x},t_{y})$ for the pair $(a_{2i},a_{2i+1})$, the total rotation radian is $\theta_{i}+\omega_{z,i}t_{z}$, where $\omega_{z,i}=b_{z}^{-\frac{2i+1}{d}}$ with a base $b_{z}$. Further details of RoPE can be found in the appendix.
Figure 2: Reconstruction results of our visual tokenizer on unseen images. The images are taken from different body parts with diverse imaging modalities and spatial properties. For each pair of images, the original image is on the left and the reconstructed image is on the right. The first row: (left) chest X-ray image from CheXpert (Irvin et al. 2019); (middle) fundus photograph from GAMMA; (right) breast ultrasound image from BUSI (Al-Dhabyani et al. 2020). The second row: (left): lung CT image from LIDC-IDRI (Armato III et al. 2011); (middle): abdomen CT image from FLARE 2022; (right): prostate MRI image from PROSTATE-MRI (Choyke et al. 2016). We clarify that the results of 3D images in the second row are obtained by processing the volume as a whole with the visual tokenizer and selecting a slice from the reconstructed volume. More examples are available in the appendix.
@@
We pre-train the ViT via visual token reconstruction with our pre-trained visual tokenizer, which is similar to BEiT (Bao et al. 2022) but the visual tokens are in soft representations. The overall method is illustrated in Figure 1. For a given 3D image $x$, it is converted into a grid of token distributions $\left\{q^{(s)}\mid s\in G_{D,H,W}\right\}$ by the pre-trained visual tokenizer. On the other hand, the patch embedding layer of ViT also divides $x$ into corresponding $D\times H\times W$ patches. We randomly mask out 55% of the patches and replace their embeddings with a learnable embedding $\bm{e}_{\mathrm{mask}}$. Transformer blocks then process the corrupted patch embeddings and output a $D\times H\times W$ grid of hidden vectors. These hidden vectors are projected to the categorical distributions $\left\{\tilde{q}^{(s)}\mid s\in G_{D,H,W}\right\}$ over $\mathcal{V}$ through a multi-layer perceptron (MLP) with a single hidden layer. The pre-training objective is minimizing the cross entropy of the reconstructed distribution relative to the distribution output by the visual tokenizer:
$L_{\mathrm{MIM}}=\frac{1}{\left\lvert\mathcal{M}\right\rvert}\sum_{s\in\mathcal{M}}\sum_{i=1}^{\left\lvert\mathcal{V}\right\rvert}q^{(s)}_{i}\ln\tilde{q}^{(s)}_{i},$ (6)
where $\mathcal{M}\subset G_{D,H,W}$ is the set of positions of all masked patches.
We intentionally choose a higher masking ratio for 3D medical images compared to previous works on 2D images (Bao et al. 2022; Fang et al. 2023a). In addition to patches within the same plane, patches from other planes may also provide useful information for reconstruction. A higher masking ratio helps balance the difficulty of the pre-training task.
Experiments
Pre-training
@@
We collect 55 public medical image datasets for both visual tokenizer and ViT pre-training, including 15 datasets of 2D images and 40 datasets of 3D images. The collected data include a wide range of image modalities (CT, MRI, X-ray, ultrasound and fundus photography) and body parts (abdomen, brain, breast, cervix, chest, fetal, fundus, head-neck, heart, pelvis, prostate and spine), with diverse spatial properties. We compare the amount and imaging modalities of data for pre-training with related works in Table 1. We pre-process the data to a unified format, including moving the depth dimension of an image to the first spatial dimension. During pre-training, we only use raw images in a self-supervised manner, and any annotations or labels from the datasets are not utilized. For each modality in each dataset, we randomly exclude one image from training data for validation. The details of all datasets for pre-training and the way of pre-processing are provided in the appendix.
@@
Table 1: The amount and imaging modalities of data used for pre-training in related and our works. Note: #3D and #2D: the number of 3D and 2D images, respectively; XR: X-ray; US: ultrasound; FP: fundus photography.
Implementation Details
For visual tokenizer, we set codebook size $|\mathcal{V}|=1024$. During pre-training, we randomly crop a $\max\{64/2^{\mathrm{DA}},1\}\times 128\times 128$ patch from an image given its DA. We adopt a batch sampling strategy that all images within a batch are selected to have the same DA, ensuring that they have the same cropped size and thus can be processed efficiently by the network in parallel. We train the visual tokenizer for 600K steps with the batch size of 48 and the peak learning rate of $10^{-4}$. For ViT pre-training, we randomly crop a $\max\{80/2^{\mathrm{DA}},1\}\times 160\times 160$ patch from an image, and use the same batch sampling strategy as the visual tokenizer training. We initialize ViT with the EVA-02-B pre-trained weights and train it for 300K steps with the batch size of 84 and the peak learning rate of $2\times 10^{-5}$. All pre-trained models are trained with AdamW optimizer (Loshchilov and Hutter 2019) and cosine learning rate scheduler with a linear warm-up. Our models are implemented with PyTorch (Paszke et al. 2019) and MONAI (Cardoso et al. 2022) and trained with NVIDIA A100 GPUs.
Results
We demonstrate the reconstruction results of our visual tokenizer on unseen validation images in Figure 2. Our universal visual tokenizer could process a wide range of medical images with heterogeneous imaging modalities and spatial properties with a unified model structure. As shown in Figure 2, images from diverse sources can be consistently reconstructed from the learned soft token representations with rich details preserved. Moreover, we conduct the following analyses to evaluate the diversity and sharpness of the learned distributions. First, following Zhang et al., we plot the codebook utilization on validation images in Figure 3. The usage percentages of all tokens are at the same order of magnitude, indicating a balanced utilization of the codebook. Second, we calculate the average value of $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ on validation images as 2.791, while $\mathrm{H}\left(p_{\mathrm{prior}}\right)=\ln|\mathcal{V}|\approx 6.931$. This result indicates that the learned distributions preserve sharpness to some extent.
Figure 3: Token utilization histogram. Every eight tokens with consecutive indexes are merged into one rectangle for better visualization in limited page width.
Classification Experiments
Dataset
MedMNIST v2 (Yang, Shi, and Ni 2021; Yang et al. 2023) is a large-scale benchmark for medical image classification on standardized MNIST-like 2D and 3D images with diverse modalities, dataset scales and tasks. It consists of 12 datasets with 708,069 2D images and 6 datasets with 9,998 3D images. The size of each image is either $28\times 28$ (2D) or $28\times 28\times 28$ (3D). We experiment on the 3D portion of MedMNIST v2, namely the Organ (Bilic, Christ et al. 2019; Xu, Zhou et al. 2019), Nodule (Armato III et al. 2011), Fracture (Jin et al. 2020a), Adrenal, Vessel (Yang et al. 2020) and Synapse datasets.
Implementation Details
Like common ViT-based models, our model performs classification tasks by feeding the hidden state of the [CLS] token into a fully-connected layer. We conduct classification experiments by fine-tuning from our and several other self-supervised pre-trained models on each dataset. Following the official split of training/validation/test sets, we train all models on the training sets for 100 epochs. The model weights from the epoch with the highest AUC score on the validation set are selected for comparison on the test set. The details on experimental settings are given in the appendix.
Methods Organ Nodule Fracture Adrenal Vessel Synapse Average
@@ Results
Table 2: Classification results on the test sets of the 3D image datasets from MedMNIST v2. The results of ResNet-50 and auto-sklearn are taken from the original paper (Yang et al. 2023). Both AUC and ACC values are shown in percentages.
@@
Table 2 compares the classification performance of our model with other models using area under the ROC curve (AUC) and accuracy (ACC) metrics. Our model consistently outperforms all other models across all datasets in terms of AUC, and achieves the highest ACC on 4 out of 6 datasets. On average, our model improves AUC by at least 4.4% and accuracy by at least 3.4% across all datasets.
Segmentation Experiments

### 24 Jan 2024 v2

@@ 4.2 SPAD-Nets
We specifically employ sum pooling for weight reduction to maintain the mean and variance of output feature values, similar to the weight inflation technique used to adapt pre-trained weights from 2D models to 3D models Zhang et al. (2022). While we have demonstrated the integration of SPAD-conv into U-Net for clarity, we note that SPAD-conv can be similarly applied to a wide range of convolution-based networks and convert them to SPAD-Nets.
Generalized Resampling
Beyond the aforementioned usages in U-Net, both downsampling and upsampling types of SPAD-conv can be generalized with a base convolution network with both kernel size and stride being $2^{k}$, where $k$ is a positive integer. Such network resamples the feature map with a ratio of $2^{k}$, and it can be used as the patch embedding in ViT or the output layer for some dense prediction tasks, such as segmentation and reconstruction. We define the degree of anisotropy (DA) of a feature map to be the minimum number of normal downsampling with a ratio of 2 performing on it to have $s_{\mathrm{slice}}<2s_{\mathrm{plane}}$ after downsampling, which yields $\mathrm{DA}=\max\left\{0,\left\lfloor\log_{2}\nicefrac{{s_{\mathrm{slice}}}}{{s_{\mathrm{plane}}}}\right\rfloor\right\}$. Since such a network can be viewed as $k$ consecutive normal resampling, for generalized downsampling, the effective number of downsampling along the depth dimension is therefore reduced to $k_{0}=\max\{k-\mathrm{DA},0\}$, with corresponding kernel size and stride reduced to $2^{k_{0}}$. The generalized upsampling can be achieved similarly by reversing the process.
4.3 Visual Tokenizer Development
We first briefly review the concept of the visual tokenizer. The visual tokenizer is implemented using a VQ-VAE van den Oord et al. (2017), a variant of VAE Kingma and Welling (2013) where the latent variables follow discrete categorical distributions over the visual codebook $\mathcal{V}$, a finite set of tokens. Specifically, an encoder learns a posterior distribution $q(z\mid x)$ to output a $D\times H\times W$ grid of categorical token distributions over $\mathcal{V}$ given an image $x$. Token samples are drawn from these categorical distributions, and converted to a $d$-dimensional embedding vector from an embedding space $\bm{e}\in\mathbb{R}^{|\mathcal{V}|\times d}$, known as the vector quantization (VQ) process. A decoder learns to reconstruct the original image given the embeddings of the drawn token samples.
We build a spatially adaptive visual tokenizer (SPAD-VT) as follows. The encoder resembles a typical convolutional backbone, but is equipped with k3s1 and downsampling types of SPAD-conv. Given an input image with voxel spacing information, it outputs a feature map with $\nicefrac{{1}}{{16}}$ of the original resolution on the in-plane dimensions and $\min\left\{\nicefrac{{2^{\mathrm{DA}}}}{{16}},1\right\}$ of the original resolution on the depth dimension. The decoder is composed of a stack of k3s1 and upsampling types of SPAD-conv. It gradually upsamples the quantized feature map output by the encoder, and finally, it outputs a reconstruction of the original image. To make the VQ process differentiable, token samples are typically drawn from the categorical distributions using Gumbel-Softmax reparameterization Maddison et al. (2017); Jang et al. (2017) and straight-through estimators Bengio et al. (2013). However, these techniques can suffer from issues of severe codebook collapse, low codebook utilization, and perturbed reconstruction process introduced by stochastic quantization Zhang et al. (2023). We propose several enhancements to make SPAD-VT produce more robust tokenization output.
@@ 4.3 Visual Tokenizer Development
Dual Prior Distribution Regularization
To mitigate the low codebook utilization issue, we adopt the prior distribution regularization (PDR) technique Zhang et al. (2023) and generalize it to the soft token representation. Let $q^{(s)}$ be the distribution of the token with spatial position $s$ output by the encoder. PDR assumes the prior that the codebook should be uniformly utilized. Thus, the averaged distribution of $q^{(s)}$ of all spatial positions should be close to the uniform distribution $p_{\mathrm{prior}}$. The objective of PDR can be represented as minimizing $D_{\mathrm{KL}}\left(\mathrm{E}\left(q^{(S)}\right)\parallel p_{\mathrm{prior}}\right)$, where $S$ is a uniformly distributed random variable of the token’s spatial position. An interpretation of PDR is available in the appendix. On the other hand, to prevent all distributions collapsing to the uniform distribution, we introduce another dual objective maximizing $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ to increase the sharpness of the learned distributions. Since $p_{\mathrm{prior}}$ is a uniform distribution, it can be shown that for any distribution $p$, $D_{\mathrm{KL}}\left(p\parallel p_{\mathrm{prior}}\right)=-\mathrm{H}\left(p\right)+\ln\left\lvert\mathcal{V}\right\rvert$, where $\mathrm{H}\left(p\right)=-\sum_{i=1}^{|\mathcal{V}|}p_{i}\ln p_{i}$ (entropy of $p$). Therefore, the dual PDR is implemented to minimize the following objective:
@@
where $\lambda_{1},\lambda_{2}$ are positive hyper-parameters. The final optimization objective for SAPD-VT is a weighted combination of the objective of VQGAN Esser et al. (2021) and $L_{\mathrm{reg}}$.
4.4 Universal Foundation Model
With previous preparation, we are now able to realize a universal foundation model for medical image analysis by building a spatially adaptive ViT (SPAD-ViT) and using the SPAD-VT to perform MIM with visual token reconstruction.
Architecture
We employ a plain ViT macro architecture that is composed of a patch embedding layer and a stack of transformer blocks Dosovitskiy et al. (2021). The patch embedding layer divides the input image into a sequence of non-overlapping patches and linearly projects each patch to an embedding vector. This process is essentially equivalent to processing the input image with a downsampling convolution that has identical kernel size and stride, and can be replaced by a downsampling type of SPAD-conv. The micro architecture of the transformer block follows EVA-02 Fang et al. (2023a), which incorporates several advancements in language models including the 2D rotary position embedding (RoPE) Su et al. (2021). We further extend 2D RoPE to work with 3D medical images, as detailed in the appendix.
@@ 4.4 Universal Foundation Model
We pre-train the SPAD-ViT via visual token reconstruction with our pre-trained SPAD-VT, which is similar to BEiT Bao et al. (2022), but the visual tokens are in soft representations. For a given 3D image $x$, it is converted into a $D\times H\times W$ grid of token distributions $\left(q^{(s)}\right)_{s\in G}$ by the pre-trained SPAD-VT, where $G$ is the set of all spatial positions on the grid. On the other hand, the patch embedding layer of ViT also divides $x$ into corresponding $D\times H\times W$ patches. We randomly mask out 55% of the patches and replace their embeddings with a learnable embedding $\bm{e}_{\mathrm{mask}}$. Transformer blocks then process the corrupted patch embeddings and output a $D\times H\times W$ grid of hidden vectors. These hidden vectors are projected to the categorical distributions $\left(\tilde{q}^{(s)}\right)_{s\in G}$ over $\mathcal{V}$ through a multi-layer perceptron (MLP) with a single hidden layer. The pre-training objective is minimizing the cross entropy of the reconstructed distribution relative to the distribution output by the SPAD-VT:
$L_{\mathrm{MIM}}=\frac{1}{\left\lvert G_{\mathrm{M}}\right\rvert}\sum_{s\in G_{\mathrm{M}}}\sum_{i=1}^{\left\lvert\mathcal{V}\right\rvert}q^{(s)}_{i}\ln\tilde{q}^{(s)}_{i},$ (3)
where $G_{\mathrm{M}}\subset G$ is the set of positions of all masked patches. We intentionally choose a higher masking ratio for 3D medical images compared to previous works on 2D images. Since patches from multiple slices may provide additional information for reconstruction, a higher masking ratio can help balance the difficulty of the pre-training task.
5 Experiments
5.1 Pre-training
@@
We collect 55 public medical image datasets for both SPAD-VT and SPAD-ViT pre-training, including 15 datasets of 2D images and 40 datasets of 3D images. The collected data include a wide range of image modalities (CT, MRI, X-ray, ultrasound and fundus photography) and body parts (abdomen, brain, breast, cervix, chest, fetal, fundus, head-neck, heart, pelvis, prostate, and spine), with diverse spatial properties. We compare the size and imaging modalities of data for pre-training with related works in Table 1. We pre-process the data to a unified format, including moving the depth dimension of an image to the first spatial dimension. During pre-training, we only use raw images in a self-supervised manner without using any annotations or labels from the datasets. For each modality in each dataset, we randomly exclude one image from training data for validation. The details of all datasets used for pre-training and the steps of pre-processing are provided in the appendix.
@@
Table 1: The size and imaging modalities of data used for pre-training in related works and this work. Note: #3D and #2D: the number of 3D and 2D images, respectively; XR: X-ray; US: ultrasound; FP: fundus photography.
Method Organ Nodule Fracture Adrenal Vessel Synapse Average
@@
Table 2: Classification results on the test sets of the 3D image datasets from MedMNIST v2. The results of ResNet-50 and auto-sklearn are taken from the original paper Yang et al. (2023). Both AUC and ACC values are shown in percentages.
Implementation Details
Our models are implemented based on MONAI Cardoso et al. (2022). For SPAD-VT, we set codebook size $|\mathcal{V}|=1024$. During pre-training, we randomly crop a $\max\{64/2^{\mathrm{DA}},1\}\times 128\times 128$ patch from an image given its DA. We adopt a batch sampling strategy in which all images within a batch are selected to have the same DA, ensuring that they have the same cropped size and thus can be processed by the network in parallel. We train SPAD-VT for 600K steps with the batch size of 48 and the peak learning rate of $10^{-4}$. For SPAD-ViT pre-training, we randomly crop a $\max\{80/2^{\mathrm{DA}},1\}\times 160\times 160$ patch from an image, and use the same batch sampling strategy as the SPAD-VT training. We initialize SPAD-ViT with the EVA-02-B pre-trained weights and train it for 300K steps with a batch size of 84 and a peak learning rate of $2\times 10^{-5}$. All pre-trained models are trained with AdamW optimizer Loshchilov and Hutter (2019).
Figure 3: Reconstruction results of SPAD-VT on unseen images. The images are taken from different body parts with diverse imaging modalities and spatial properties. For each pair of images, the original image is on the left and the reconstructed image is on the right. The first row: (left) chest X-ray image from CheXpert Irvin et al. (2019); (middle) fundus photograph from GAMMA; (right) breast ultrasound image from BUSI Al-Dhabyani et al. (2020). The second row: (left): lung CT image from LIDC-IDRI Armato III et al. (2011); (middle): abdomen CT image from FLARE 2022; (right): prostate MRI image from PROSTATE-MRI Choyke et al. (2016). More examples are available in the appendix.
Figure 4: Token utilization histogram. Every eight tokens with consecutive indexes are merged into one rectangle for better visualization within the limited page width.
@@ Results
We demonstrate the reconstruction results of SPAD-VT on unseen validation images in Figure 3. SPAD-VT could process a wide range of medical images with heterogeneous imaging modalities and spatial properties using a unified model structure. As shown in Figure 3, images from diverse sources can be consistently reconstructed from the learned soft token representations with rich details preserved. Moreover, we conduct the following analyses to evaluate the diversity and sharpness of the learned distributions. First, we plot the codebook utilization on validation images in Figure 4. The usage percentages of all tokens are in the same order of magnitude, indicating a balanced utilization of the codebook. Second, we calculate the average value of $\mathrm{E}\left(D_{\mathrm{KL}}\left(q^{(S)}\parallel p_{\mathrm{prior}}\right)\right)$ on validation images as 2.791, while $\mathrm{H}\left(p_{\mathrm{prior}}\right)=\ln|\mathcal{V}|\approx 6.931$. This result indicates that the learned distributions preserve sharpness to some extent.
5.2 Classification Experiments